# CoAct-1: Computer-using Agents with Coding as Actions

### v2

@@ Abstract
Autonomous agents that operate computers via Graphical User Interfaces (GUIs) often struggle with efficiency and reliability on complex, long-horizon tasks. While augmenting these agents with planners can improve task decomposition, they remain constrained by the inherent limitations of performing all actions through GUI manipulation, leading to brittleness and inefficiency. In this work, we introduce a more robust and flexible paradigm: enabling agents to use coding as an enhanced action. We present CoAct-1, a novel multi-agent system that synergistically combines GUI-based control with direct programmatic execution. CoAct-1 features an Orchestrator that dynamically delegates subtasks to either a conventional GUI Operator or a specialized Programmer agent, which can write and execute Python or Bash scripts. This hybrid approach allows the agent to bypass inefficient GUI action sequences for tasks like file management and data processing, while still leveraging visual interaction when necessary. We evaluate our system on the challenging OSWorld benchmark, where CoAct-1 achieves a new state-of-the-art success rate of 60.76%, significantly outperforming prior methods. Furthermore, our approach dramatically improves efficiency, reducing the average number of steps required to complete a task to just 10.15, compared to 15 for leading GUI agents. Our results demonstrate that integrating coding as a core action provides a more powerful, efficient, and scalable path toward generalized computer automation.
[Uncaptioned image] Website: https://linxins.net/coact/
@@ 1 Introduction
Recent advancements in computer-using agents have primarily focused on operating through Graphical User Interfaces (GUIs). While these GUI agents, powered by vision-language (action) models (Li et al., 2023, Deepmind, 2025a, b, OpenAI, 2024, 2025a, Qin et al., 2025, Xie et al., 2025a, Yang et al., 2025), have demonstrated the ability to perform a variety of tasks, they often struggle with long-horizon planning and interactions in environments with dense GUI elements. For example, workflows common in office productivity suites or involving direct OS-level operations—such as locating a specific table within a multi-sheet spreadsheet, filtering it based on complex criteria, copying the results, and then saving it as a new CSV file—often involve a long and intricate sequence of precise GUI manipulations. Similarly, tasks like finding all image files in a nested directory structure, resizing them to specific dimensions, and then compressing the entire directory into a single archive are brittle and inefficient when solved via GUI actions such as clicking and dragging. In these scenarios, existing agents frequently struggle with visual grounding ambiguity (e.g., distinguishing between visually similar icons or menu items) and the accumulated probability of making any single error over the long horizon. A single mis-click or misunderstood UI element can derail the entire task.
@@
To mitigate these challenges, a prominent line of research has focused on augmenting GUI agents with dedicated high-level planners. Approaches such as GTA-1 (Yang et al., 2025) and other modular systems (Yang et al., 2024, Xu et al., 2024, Agashe et al., 2024, 2025) employ powerful language models like OpenAI o3 (OpenAI, 2025b) to decompose a user’s high-level goal into a sequence of more manageable subtasks. This hierarchical decomposition can improve performance on complex, multi-step problems by providing a structured plan. However, this paradigm does not fundamentally address the inefficiency and brittleness inherent in relying exclusively on GUI-based actions for execution. The agent must still navigate menus, click buttons, and type into fields, even for operations that could be accomplished more directly and reliably through programmatic means. This leaves the system susceptible to planning uncertainty, visual perception errors, and the integration challenges between high-level planning and low-level action generation.
@@
In this work, we argue for and instantiate a more flexible and powerful action space. We propose a hybrid approach that combines the intuitive, human-like strengths of GUI manipulation with the precision, reliability, and efficiency of direct system interaction through code. We introduce CoAct-1 (Computer-using Agent with Coding as Actions), a novel multi-agent system that is architected around three specialized agents: Orchestrator, Programmer, and GUI Operator. A high-level Orchestrator serves as the central planner, decomposing the user’s goal and assessing the nature of each subtask. Based on this analysis, it assigns the task to one of two distinct execution agents: a Programmer agent, which writes and executes Python or Bash scripts for backend operations like file management, data processing, or environment configuration; or a GUI Operator, a VLM-based agent that performs frontend actions like clicking buttons and navigating visual interfaces. This dynamic delegation allows CoAct-1 to strategically bypass inefficient GUI sequences in favor of robust, single-shot code execution where appropriate, while still leveraging visual interaction for tasks where it is indispensable.
@@
Our experimental analysis provides strong evidence for the advantages of this hybrid design. On the comprehensive OSWorld benchmark, CoAct-1 establishes a new state-of-the-art, achieving an overall success rate of 60.76%. This marks a significant improvement over leading baselines like GTA-1 (53.10%). The performance gains are most pronounced in categories where programmatic control is highly advantageous. For instance, in Calc (70.21%), multi-application (47.88%), and VS Code (78.26%) tasks, the Programmer agent’s ability to execute precise scripts leads to substantial gains over the strongest GUI-only methods. Beyond improving success rates, our dual-modality approach dramatically enhances operational efficiency. By replacing long, error-prone click sequences with concise code, CoAct-1 solves tasks in an average of just 10.15 steps, a stark contrast to the 15 steps required by agents like GTA-1. This efficiency underscores the potential of our approach to pave a more robust and scalable path toward generalized computer automation.
@@ 2 Related Work
Screen parsing and visual grounding A first line of work focuses on perceiving and grounding GUI elements directly from pixels, without relying on DOM or accessibility hooks. OmniParser learns screen-parsing primitives for pure vision–based understanding (Lu et al., 2024). On the grounding side, SeeClick (instruction-to-target grounding), Aria-UI (instruction grounding over GUIs), and UGround (universal GUI grounding) map language to actionable screen locations (Cheng et al., 2024, Yang et al., 2024, Gou et al., 2024). OS-Atlas trains a foundation action model to generalize across diverse interfaces (Wu et al., 2024). Dedicated grounding evaluations such as ScreenSpot-Pro further benchmark grounding under professional, high-resolution settings (Li et al., 2025).
Modular planner–grounder agents A second strand explicitly separates what to do from where/how to act on screen: a language planner proposes subgoals while a visual model grounds each step. Representative systems include SeeClick and OS-Atlas (Cheng et al., 2024, Wu et al., 2024). GTA-1 strengthens this two-stage paradigm via test-time scaling: sampling multiple candidate actions and using an MLLM judge to select among them, improving robustness on high-resolution, cluttered UIs (Yang et al., 2025). Other related open frameworks such as Agent-S / Agent-S2 and AutoGen provide reusable infrastructures for multi-agent orchestration and tool calling (Agashe et al., 2024, 2025, Song et al., 2025, Zhang et al., 2024, Wu et al., 2023, Zhang et al., 2023, 2025b).
Native end-to-end GUI agents A third thread trains native agents that unify perception, reasoning, and action in a single model. UI-TARS exemplifies this approach with a unified action space for mouse/keyboard operations across apps, eschewing hand-crafted controllers (Qin et al., 2025). AGUVIS pushes toward unified, pure-vision GUI agents that generalize across interfaces (Xu et al., 2024).
Hybrid agentic frameworks Beyond GUI-only interaction, several agentic systems compose tools and APIs on the fly to extend capabilities at run time. Examples include UFO-2 (Zhang et al., 2025a), PyVision (Zhao et al., 2025), and ALITA (Qiu et al., 2025)—which, while not restricted to GUI/CUA, share the principle of dynamically constructing and invoking tools.
@@ 3 Computer-using Agent with Coding as Actions
Figure 1: Multi-agent system design for our CoAct-1. This multi-agent system includes a programmer who can interact with the operating system through multi-round coding.
In this work, we introduce a new system-interactive action: coding, to replace redundant and brittle GUI actions. Specifically, we design a multi-agent system called CoAct-1 based on the GUI agent with a planner approach, introducing a new agent called programmer who can interact with the operating system by coding. An orchestrator is designed to determine whether to assign the programmer or the GUI operator to complete a subtask. The overall framework is illustrated in Figure 1.
@@ 3.1 Multi-agent System Design for Computer Use
Our multi-agent system comprises three agents: the Orchestrator, Programmer, and GUI Operator, each of which will establish a conversation.
Orchestrator. The orchestrator is responsible for decomposing user tasks and dynamic planning. The orchestrator’s step-wise action is based on all previous observations. Orchestrator cannot interact with the OS directly. Therefore, in each step, the orchestrator is required to assign a subtask to either a programmer or a GUI operator. After the programmer and the GUI operator complete their subtask, the operator will receive a summarization of their task-solving process and a screenshot that reflects the current system state. If the orchestrator determines that the task is complete, it will return a termination signal to end the conversation.
Programmer. Once the orchestrator assigns a subtask to the programmer, we will establish a new conversation between the coding agent and code interpreter, enabling the coding agent to reflect the generated code according to the environment feedback in multiple rounds. The coding agent is required to solve the assigned subtask by programming with Python or Bash scripts, which is the two most frequently used languages when operating the OS. When the code interpreter successfully retrieves the code from the coding agent’s reply, it sends the code to the operating system (which can be a remote virtual machine). Then it returns the execution result to the conversation along with a screenshot. Notably, we use a language model to serve as a coding agent with limited ability to read and understand sequential images. Therefore, we require the orchestrator to provide sufficient environment information to the programmer, such as the opened windows, browser tabs, or file paths inferred from the user’s instructions.
GUI operator. The GUI agent is a vision-language action model, which can take a sequence of images along with the text instruction as input and generate a (sequence of) GUI action(s). When the GUI operator receives a subtask from the orchestrator, it will establish a conversation between a GUI agent and a GUI action interpreter. The GUI action includes mouse movement, mouse clicks, keyboard hotkeys, and keyboard typing. The GUI interpreter is a non-language model, proxy agent responsible for interpreting the language model’s actions as system operations, and returning the resulting screenshot to the GUI agent. Once the GUI agent decides to complete the task, it will output a message along with a termination signal to end the conversation. This message includes the information that the orchestrator requires the GUI agent to collect, such as a file path, and will be returned to the orchestrator along with a screenshot.
@@ 3.2 Workflow and Memory Design
For each agent, we use the instance conversation history as their memory. As illustrated in Figure 2, when the Programmer completes an assigned subtask, a language model will summarize the conversation between the coding agent and code interpreter, returning a summary and a screenshot to the Orchestrator as part of the Orchestrator’s memory. On the other hand, the GUI Operator will return a message that includes the required information from the Orchestrator, which is also stored as part of the Orchestrator’s memory. The Programmer, Orchestrator, and GUI Operator do not share their conversation history. Once the subtask is completed, the system will clean the instance conversation history of the Programmer and the GUI Operator, allowing them to focus solely on the current assigned subtask.
@@ 4 Experiments
4.1 Benchmark Dataset
We test the CoAct-1 and other baselines on the OSWorld benchmark (Xie et al., 2024). OSWorld is a scalable real-computer testbed that exposes an OS like Linux to an agent through both pixel streams and an OS shell interface. The benchmark currently comprises 369 tasks that span common productivity tools, IDEs, browsers, file managers, and multi-application workflows, thereby stressing both vision–language grounding and long-horizon planning in heterogeneous GUI environments. Each task in OSWorld ships with (i) a deterministic VM snapshot capturing the initial desktop state, (ii) a natural-language goal that mirrors how end-users phrase requests (e.g., “resize the image to $512\times 512$ and export as PNG”), and (iii) a rule-based evaluator. Tasks cover atomic operations (single-app editing, file I/O) as well as cross-app pipelines, such as “export this spreadsheet, compress it, then e-mail the archive,” providing a realistic spectrum of difficulty.

### v3

# CoAct-1: Computer-using Multi-agent System with Coding Actions

@@ Abstract
Autonomous agents that operate computers via Graphical User Interfaces (GUIs) often struggle with efficiency and reliability on complex, long-horizon tasks. While augmenting these agents with planners can improve task decomposition, they remain constrained by the inherent limitations of performing all actions through GUI manipulation, leading to brittleness and inefficiency. In this work, we introduce a more robust and flexible paradigm: enabling agents to use coding as an enhanced action. We present CoAct-1, a novel multi-agent system that synergistically combines GUI-based control with direct programmatic execution. CoAct-1 features an Orchestrator that dynamically delegates subtasks to either a conventional GUI Operator or a specialized Programmer agent, which can write and execute Python or Bash scripts. This hybrid approach allows the agent to bypass inefficient GUI action sequences for tasks like file management and data processing, while still utilizing visual interaction when necessary. We evaluate our system on the challenging OSWorld and WindowsAgentArena benchmark, where CoAct-1 achieves a new state-of-the-art success rate of 60.8% on OSWorld and 52.5% on WindowsAgentArena, significantly outperforming prior methods. Furthermore, our approach dramatically improves efficiency, reducing the average number of steps required to complete a task to just 10.15 on OSWorld, compared to 15 for leading GUI agents. Our results demonstrate that integrating coding as a core action provides a more powerful, efficient, and scalable path toward generalized computer automation.
@@ 1 Introduction
Recent advancements in computer-using agents have primarily focused on operating through Graphical User Interfaces (GUIs). While these GUI agents, powered by vision-language (action) models (Li et al., 2023; Deepmind, 2025b; a; OpenAI, 2024; 2025a; Qin et al., 2025; Xie et al., 2025b; Yang et al., 2025), have demonstrated the ability to perform a variety of tasks, they often struggle with long-horizon planning and interactions in environments with dense GUI elements. For example, routine tasks in Office productivity software often involve a long and intricate sequence of precise GUI operations, such as locating a specific table within a multi-sheet spreadsheet, filtering it based on complex criteria, copying the results, and saving them as a new CSV file. Similarly, tasks like finding all image files in a nested directory structure, resizing them to specific dimensions, and then compressing the entire directory into a single archive are brittle and inefficient when solved via GUI actions such as clicking and dragging. In these scenarios, existing agents often struggle with visual grounding ambiguity (e.g., mistaking visually similar icons or menu items) and the cumulative probability of errors over long-term interactions. A single mis-click or misunderstood UI element can derail the entire task.
@@
To address these challenges, a prominent line of research has focused on augmenting GUI agents with dedicated high-level planners. Approaches such as GTA-1 (Yang et al., 2025) and other modular systems (Yang et al., 2024; Xu et al., 2024; Agashe et al., 2024; 2025) utilize powerful language models like OpenAI o3 (OpenAI, 2025b) to decompose a user’s high-level goal into a sequence of more manageable subtasks. This hierarchical decomposition can improve performance on complex, multi-step problems by providing a structured plan. However, this paradigm does not fundamentally address the inefficiency and brittleness associated with exclusive reliance on GUI-based execution. Even with the high-level planning, the agent still needs to navigate menus, click buttons, and type into fields, even for operations that could be accomplished more directly and reliably through programmatic means. This leaves the system susceptible to planning uncertainty, visual perception errors, and the integration challenges between high-level planning and low-level action generation.
@@
In this work, we advocate for and instantiate a more flexible and powerful action space. We propose a hybrid approach that combines the intuitive, human-like strengths of GUI manipulation with the precision, reliability, and efficiency of direct system interaction through code. We introduce CoAct-1 (Computer-using Multi-agent System with Coding Actions), a novel multi-agent system composed of three specialized agents: Orchestrator, Programmer, and GUI Operator. A high-level Orchestrator serves as the central planner, decomposing the user’s goal and determining the appropriate modality for each subtask. Based on this analysis, it assigns the task to one of two distinct execution agents: a Programmer agent, which writes and executes Python or Bash scripts for backend operations like file management, data processing, or environment configuration; or a GUI Operator, a VLM-based agent that performs frontend actions like clicking buttons and navigating visual interfaces. This dynamic delegation allows CoAct-1 to strategically bypass inefficient GUI sequences in favor of robust, single-shot code execution when appropriate, while still leveraging visual interaction for tasks.
@@
Our experimental analysis provides strong evidence for the advantages of this hybrid design. On the OSWorld and WindowsAgentArena benchmark, CoAct-1 establishes a new state-of-the-art, achieving an overall success rate of 60.76% and 52.50%, respectively. This marks a significant improvement over leading baselines like Agent S2.5 (55.98%) on OSWorld. The performance gains are most pronounced in categories where programmatic control is highly advantageous. For instance, in Calc (70.21%), multi-application (47.88%), and VS Code (78.26%) tasks, our Programmer’s ability in executing precise scripts leads to substantial gains over the strongest GUI-only methods. Beyond improving success rates, our dual-modality approach dramatically enhances operational efficiency. By replacing long, error-prone click sequences with concise code, CoAct-1 solves tasks in an average of just 10.15 steps on OSWorld, a stark contrast to the 15 steps required by agents like GTA-1. This efficiency underscores the potential of our approach to pave a more robust and scalable path toward generalized computer automation.
@@ 2 Related Work
Screen parsing and visual grounding A first line of work focuses on perceiving and grounding GUI elements directly from pixels, without relying on DOM or accessibility hooks. OmniParser learns screen-parsing primitives for pure vision–based understanding (Lu et al., 2024). On the grounding side, SeeClick (instruction-to-target grounding), Aria-UI (instruction grounding over GUIs), and UGround (universal GUI grounding) map language to actionable screen locations (Cheng et al., 2024; Yang et al., 2024; Gou et al., 2024). OS-Atlas trains a foundation action model to generalize across diverse interfaces (Wu et al., 2024). Dedicated grounding evaluations such as ScreenSpot-Pro further benchmark grounding under professional, high-resolution settings (Li et al., 2025).
Native end-to-end GUI agents A third thread trains native agents that unify perception, reasoning, and action in a single model. UI-TARS and OpenCUA exemplifies this approach with a unified action space for mouse/keyboard operations across apps, eschewing hand-crafted controllers (Qin et al., 2025; Wang et al., 2025a; b). AGUVIS pushes toward unified, pure-vision GUI agents that generalize across interfaces (Xu et al., 2024).
Modular planner–grounder agents A second strand explicitly separates what to do from where/how to act on screen: a language planner proposes subgoals while a visual model grounds each step. Representative systems include SeeClick and OS-Atlas (Cheng et al., 2024; Wu et al., 2024). GTA-1 strengthens this two-stage paradigm via test-time scaling: sampling multiple candidate actions and using an MLLM judge to select among them, improving robustness on high-resolution, cluttered UIs (Yang et al., 2025). Other related open frameworks such as Agent-S / Agent-S2 and AutoGen provide reusable infrastructures for multi-agent orchestration and tool calling (Agashe et al., 2024; 2025; Song et al., 2025a; Zhang et al., 2024; Wu et al., 2023; Zhang et al., 2023; 2025b).
Hybrid agentic frameworks Beyond GUI-only interaction, several agentic systems compose tools and APIs on the fly to extend capabilities at run time. Examples include UFO-2 (Zhang et al., 2025a), PyVision (Zhao et al., 2025), BeyondBrowsing (Song et al., 2025b) and ALITA (Qiu et al., 2025), which, while not restricted to GUI/CUA, share the principle of dynamically constructing and invoking tools.
@@ 3 Computer-using Agent with Coding as Actions
Figure 1: Multi-agent system design for our CoAct-1. This multi-agent system includes a Programmer that can interact with the operating system through multi-round coding. This multi-agent system includes an Orchestrator, which serves as the high-level planner that decomposes goals and delegates subtasks to the appropriate execution agent, a Programmer, which interacts with the operating system through multi-round coding, and a GUI Operator, which leverages vision-language capabilities to perform visual interface actions.
In this work, we introduce a new system-interactive action: coding, to replace part of the redundant and brittle GUI actions. Unlike summarizing APIs or SDKs from each application or website, we focus on enabling agents to perform free-form coding to solve computer-use problems guided by a strong language model. Specifically, we design a multi-agent system, CoAct-1, that introduces a new agent, Programmer, capable of interacting with the OS through a coding-observation circle. An Orchestrator serves as the high-level controller, determining whether to assign the subtask to the Programmer or the GUI Operator. The overall framework is illustrated in Figure 1.
3.1 Problem Definition
We formalize the problem of general computer control as an interactive decision-making process. At each timestep $t$, the agent observes the computer environment (primarily consists of a screenshot) $o_{t}\in{\mathcal{O}}$, and takes an action $a_{t}\in{\mathcal{A}}$ according to a policy $\pi(a_{t}|H_{t},G)$. Here, $H_{t}=(o_{1},a_{1},...,o_{t-1},a_{t-1},o_{t})$ represents the historical context, and $G$ is the user’s high-level goal provided in natural language. Learning an effective policy is particularly challenging when the action space $\mathcal{A}$} is restricted to low-level GUI operations. Complex tasks, such as managing nested files or processing spreadsheet data, can require long and intricate sequences of GUI actions. This makes the process inefficient and highly susceptible to error propagation, where a single mis-check can derail the entire task.
To address this limitation, we introduce a hybrid action space that integrates direct programmatic control. We augment it to the standard GUI action space, denoted as ${\mathcal{A}}={\mathcal{A}}_{\text{GUI}}\cup{\mathcal{A}}_{\text{Code}}$. An action $a_{t}\in{\mathcal{A}}_{\text{GUI}}$ involves the direct manipulation of the graphical interface (e.g., mouse clicks, keyboard typing). In contrast, an action $a_{t}\in{\mathcal{A}}_{\text{Code}}$ consists of a Python or Bash script that interacts directly with the operating system’s backend. This allows the agent to perform complex operations like file manipulation or data processing in a single, robust step, effectively bypassing brittle and inefficient GUI sequences. In CoAct-1, the policy $\pi$ is implemented hierarchically. A high-level Orchestrator acts as a meta-policy, $\pi_{\text{orch}}$, which analyzes the current subtask and delegates it to one of two specialized executor policies: a GUI Operator that implements $\pi_{\text{GUI}}$ for actions in ${\mathcal{A}}_{\text{GUI}}$, or a Programmer that implements $\pi_{\text{Code}}$ for actions in ${\mathcal{A}}_{\text{Code}}$.
@@ 3.2 Multi-agent System Design for Computer Use
Our multi-agent system is the architectural instantiation of the hierarchical policy $\pi$ outlined in the problem definition. It comprises three specialized agents—the Orchestrator, Programmer, and GUI Operator that collaboratively generate the action trajectory $\tau$ to solve the user’s goal $G$. Each agent establishes a dedicated conversation to perform its role.
Orchestrator. The Orchestrator embodies the high-level meta-policy, $\pi_{\text{Orch}}$. It is responsible for task decomposition and dynamic planning based on the full history of observations $H_{t}$ and the overall goal $G$. The Orchestrator does not interact directly with the OS. Instead, its primary function is to select the best specialized sub-policy, $\pi_{\text{Code}}$ or $\pi_{\text{GUI}}$, to execute the current subtask. Upon completion, the Orchestrator receives a summary of the execution process and a new observation $o_{t+1}$ (a screenshot along with a summary reflecting the current system state) to inform its next decision. If it determines that the overall goal $G$ is met, it outputs a termination signal.
Programmer. The Programmer implements the specialized policy $\pi_{\text{Code}}$, responsible for generating actions $a_{t}\in{\mathcal{A}}_{\text{Code}}$. Upon receiving a subtask from the Orchestrator, it initiates a multi-round conversation with a code interpreter. It generates Python or Bash scripts, which are executed by the interpreter. The feedback, consisting of the code execution results, allows the Programmer to reflect and refine their code until the subtask is solved. The Orchestrator provides the Programmer with sufficient context, such as file paths or window information inferred from $H_{t}$, to ground its code generation.
GUI operator. The GUI agent is a vision-language action model that implements the GUI-based policy $\pi_{\text{GUI}}$ for generating actions $a_{t}\in{\mathcal{A}}_{\text{GUI}}$. Similar to the Programmer, the GUI Operator engages in a multi-round interactive loop to complete its assigned subtasks. In each step of this "perception-action" loop, the agent takes the current screenshot and the subtask instruction as input to generate a single GUI action (e.g., a mouse click or keyboard input). A GUI action interpreter executes this action on the OS, which in turn provides a new screenshot as observational feedback. The GUI Operator uses this new visual information to decide on its subsequent action, continuing this cycle until its subtask is complete.
@@ 3.3 Workflow and Memory Design
The hierarchical policy implemented by CoAct-1 necessitates a structured workflow and memory system to manage the flow of information between agents. This design ensures that each agent has the necessary context without being overwhelmed by irrelevant details from other parts of the task. The overall workflow is illustrated in Figure 2.
Workflow
The workflow begins when the Orchestrator, acting as the meta-policy $\pi_{\text{Orch}}$, delegates a subtask to an appropriate executor agent: either the Programmer or the GUI Operator. The selected agent then engages in its own multi-round interactive loop to solve the subtask, generating a detailed conversational history of its actions and the environment’s responses. Upon completing the subtask, this detailed history is processed by a dedicated summarizer model. This model condenses the entire interaction into a concise summary that captures the key actions taken and the final outcome. This summary, along with the final screenshot representing the new state of the environment, is then returned to the Orchestrator. This handoff mechanism provides the Orchestrator with a condensed, high-level update to its overall task history $H_{t}$.
Memory design for task decomposition
CoAct-1 employs a hierarchical and isolated memory structure:
  • Orchestrator Memory: The Orchestrator maintains the long-term, primary memory, which corresponds to the historical context $H_{t}$ from our problem definition. It consists of the initial user goal $G$ and the sequence of summaries and screenshots received from the executor agents after each completed subtask. This aggregated history provides the context for all high-level planning decisions.
  • Executor Memory: The Programmer and GUI Operator each maintain a short-term, working memory that is active only for the duration of their assigned subtask. This memory contains the "instance conversation history" of their multi-round interaction with the OS.
To ensure modularity and focus, these memories are isolated; the agents do not share their conversational histories directly. Furthermore, once an executor completes its subtask and reports back to the Orchestrator, its working memory is cleared. This reset mechanism is critical, as it allows the executor agents to focus entirely on the context of the new subtask they receive without being influenced by prior, irrelevant interactions.
@@ 4 Experiments
4.1 Benchmark Datasets
We evaluate CoAct-1 on OSWorld (Xie et al., 2024) and WindowsAgentArena (Bonatti et al., 2024). Both are scalable real‑computer testbed that exposes an OS (Windows or Ubuntu) to an agent through pixel streams and an OS shell interface. OSWorld comprises 369 tasks, while WindowsAgentArena comprises 154 tasks. These task span common productivity tools, IDEs, browsers, file managers, and multi-application workflows, thereby challenging both vision–language grounding and long-horizon planning in heterogeneous GUI environments.